OpenStudy (anonymous):

A firecracker shoots up from a hill 160 feet high with an initial speed of 90 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (anonymous):

this is beyond hard!

OpenStudy (anonymous):

thx real big help XD

OpenStudy (anonymous):

do get awarded for this, lol?

OpenStudy (anonymous):

I got this! Lol! 1 sec!

OpenStudy (anonymous):

k thx jon can you show how you do it? thanks

OpenStudy (agent0smith):

First plug in the speed v they give you, 90, and i'm guessing the 160 is the s \[\Large H(t) = −16t^2 + 90t + 160\] The height when it reaches the ground will be zero, which means H is 0: \[\Large 0 = −16t^2 + 90t + 160\] Now you have to solve this, either by factoring or using the quadratic formula. To make it a little easier, you can first divide everything by 2 to get \[\Large 0 = −8t^2 + 45t + 80\]

OpenStudy (anonymous):

0=−16t2+105t+160 A standard quadratic is of the form ax2+bx+c=0 its solution is given as x=−b±b2−4ac−−−−−−−√2a here b=105 a=-16 c=160 Let's plugin the values x=−(105)±(105)2−4×(−16)×(160)−−−−−−−−−−−−−−−−−−−−−√2×(−16) we get x=−(105)±21265−−−−−√2×(−16) x=−(105)±145.8252×(−16) if we take + sign then t= -ve ( time can't be negative) with - sign x=−250.825−32

OpenStudy (agent0smith):

Quadratic formula is probably easier.. do you know how to use it? \[\Large t = \frac{ -b \pm \sqrt {b^2 - 4 ac} }{ 2a }\]

OpenStudy (agent0smith):

To get a, b and c, compare your function: \[\Large 0=−8t^2+45t+80\]to the standard \[\Large 0=at^2+bt+c\]

OpenStudy (agent0smith):

@Jonathan1997 where did 105 come from...? Your solution isn't very easy to read...

OpenStudy (anonymous):

in the quadractic formula on the inside of it you would have 16 -14,400 and then you get -14,416 right? but i thought things inside the radican cant be negative?

OpenStudy (anonymous):

Okay, so you have s = 160, v = 105 so h(t) = -16t^2 + 105t + 160, use quadratic equation to find roots of this equation, the one with a positive t value will be your answer.

OpenStudy (agent0smith):

plug in a, b and c... from above, a=-8, b=45, c=80\[\Large t = \frac{ -45 \pm \sqrt {45^2 - 4 *(-8)*80} }{ 2(-8) }\]now simplify that, carefully, pay attention to negatives.

OpenStudy (agent0smith):

@Jonathan1997 v is not 105.

OpenStudy (anonymous):

Okay, well I will just let agenta.s.s. help ya! GOOD LUCK!!! You will need it @Allyt

OpenStudy (anonymous):

okayy . @agent0smith it would be 7.04375 but it would be 7 secounds right?

OpenStudy (agent0smith):

Correct.

OpenStudy (anonymous):

*cough* *cough* sorry @agent0smith